# A Path to Constraints on Common Envelope Ejection in Massive Binaries: Full Evolutionary Reconstruction of Three Black Hole X-ray Binaries

### 12 Apr 2026 v1

@@ Abstract
The massive binary common envelope (CE) phase plays a pivotal role in the formation of close black hole/neutron star (BH/NS) binaries, yet significant uncertainties remain in our understanding of this process. In this study, we aim to constrain the massive binary CE phase by systematically reconstructing three observed BH X-ray binaries (BHXBs): GRO J1655-40, SAX J1819.3-2525, and 4U 1543-47. Through comprehensive binary evolution simulations and parametric supernova (SN) modeling, we establish lower limits for the CE efficiency parameters under different energy considerations within the standard energy formalism. Specifically, we derive minimum values for three cases: $\alpha_{\rm 0.5U}$ and $\alpha_{\rm U}$ representing CE efficiencies with half and all of the internal energy contributing to the envelope ejection, respectively, and $\alpha_{\rm H}$ accounting for the envelope’s enthalpy. Our analysis reveals that the self-consistent formation of these three BHXBs requires CE efficiency parameters satisfying: $\alpha_{\rm 0.5U}\gtrsim 6.7$, $\alpha_{\rm U}\gtrsim 4.2$ and $\alpha_{\rm H}\gtrsim 1.7$. Notably, we find no viable solutions with CE efficiency values below unity, even when considering the most extreme scenarios in which the envelope binding energy is significantly reduced through enthalpy inclusion. Our results strongly imply that either additional energy sources are required, or the formalism itself must be revised. Furthermore, we quantitatively assess the impact of BH natal kicks on our results. A key finding is that 4U 1543-47’s formation requires substantial natal kicks ($\gtrsim 50\;\rm km/s$), as lower kick velocities are incompatible with isolated binary evolution.
@@
software: MESA (v12115; Paxton et al. 2011, 2013, 2015, 2018, 2019)
@@ I Introduction
Although significant effort has been devoted to modeling CE evolution through numerical simulations spanning 1D to 3D approaches (e.g., Soker, 2016; Ohlmann et al., 2016; Ivanova and Nandez, 2016; Kruckow et al., 2016; Iaconi et al., 2017; Shiber and Soker, 2018; Chamandy et al., 2018; Fragos et al., 2019; Shiber et al., 2019; Law-Smith et al., 2020; Grichener et al., 2021; Moreno et al., 2022; Lau et al., 2022; Bronner et al., 2024; Gagnier and Pejcha, 2025; Vetter et al., 2025), the astrophysical community has largely adopted a phenomenological description based on energy conservation as a practical framework for modeling the CE ejection (Paczynski, 1976; Webbink, 1984; Livio and Soker, 1988). In this energy formalism [^1] In the remainder of this paper, the CE energy formalism refers to the standard version used in binary population synthesis studies. It therefore does not consider additional processes or energy sources that may operate during the CE phase, including but not limited to jets (e.g., Soker 2004, 2015, 2016; Shiber et al. 2019), magnetic fields (e.g., Ohlmann et al. 2016; García-Segura et al. 2020) and nuclear energy Podsiadlowski et al. 2010., the system’s orbital energy release is assumed to overcome the envelope’s binding energy. This approach offers significant advantages: its simplicity and computational efficiency make it particularly suitable for population synthesis studies (see recent reviews by Han et al. 2020; Chen et al. 2024).
@@ II.2 Formation scenario of BHXBs
Figure 1: The typical evolutionary scenario for BH I/LMXBs. Abbreviations are as follows: ZAMS–zero-age main sequence, MT–mass transfer, CE–common envelope, He–helium, MS–main sequence, SN–supernova, BH–black hole, I/LMXB–intermediate-/low-mass X-ray binary. At stage 4, the dashed contour denotes the Roche lobe of the secondary star. Note that the successful ejection of the common envelope requires the secondary’s radius to be within its Roche lobe.
@@ II.3 The technical route
where $P$ and $\rho$ represent the pressure and density of each mass shell. The quantity $U+P/\rho$ is generally known as enthalpyenthalpy $(H)$. It should be noted that enthalpy consideration does not introduce new energy sources, but rather redistributes the deposited heat during the spiral-in phase (Ivanova and Chaichenets, 2011; Ivanova et al., 2013). Following Ivanova (2011), we define the core-envelope boundary as the point with the maximum compression $P/\rho$ inside the hydrogen burning shell (characterized by the peak values of the local sonic velocity; see also Marchant et al. 2021; Vigna-Gómez et al. 2022).
Figure 2: Binding energy calculation example for a $34\,M_{\odot}$ primary star. Panel (a): The evolutionary track in the HR diagram with the red cross indicating the specific evolutionary stage analyzed. The thick green line indicates that the star contains a helium core. Panel (b): Radial profiles of hydrogen mass fraction ($X_{H}$) and the sonic velocity $C_{\rm s}\equiv\sqrt{P/\rho}$ at this stage, where open circles mark the core-envelope boundary. Panel (c): The evolution of the primary envelope energy components (gravitational energy $E_{\rm gr}$, internal energy $E_{\rm U}$, and enthalpy term $E_{\rm H}$), as functions of stellar radius at each specific evolutionary stage (covering the post-MS stages). Panel (d): Similar to panel (c), but for the integrated envelope binding energy vs. radius for the three cases.
@@
From Stage 5 to Stage 4, the orbit evolution is governed by mass loss through the stripped He star’s stellar wind, which we consider as a fast (several $1000\;\rm km/s$; Vink 2022) and isotropic spherically-symmetric wind that does not interact with the companion star. The assumptions are well-justified for Wolf-Rayet stars where the wind velocity significantly exceeds the typical orbital velocity ($\sim 100\;\rm km/s$), making wind accretion negligible (El Mellah et al., 2019). The change in binary separation due to the wind mass loss then is given by (Postnov and Yungelson, 2014)
@@
For each set of binary parameters at Stage 6 that successfully reproduces the observed BHXB properties, we can systematically reconstruct the corresponding post-CE binary configuration at Stage 4. Using Equation (1) with the calculated binding energy $E_{\rm bind}$, we then determine the required CE efficiency parameter $\alpha_{\rm CE}$ for each viable solution.
@@ II.5 The input parameters
We perform all stellar evolution calculations using the state-of-the-art stellar evolutionary code Modules for Experiments in Stellar Astrophysics (MESA; version 12115; Paxton et al. 2011, 2013, 2015, 2018, 2019), adopting solar metallicity (Z=0.0142) with initial He fraction Y=0.2703 (Asplund et al., 2009) for non-rotating massive star models (as the strong stellar winds in massive stars cause a marked spin-down; e.g., Nathaniel et al. 2025). Our simulations incorporate Type II opacities (Iglesias and Rogers, 1996) and the approx21_\_cr60_\_plus_\_co56.net nuclear reaction network (Ferguson et al., 2005), with convection parameters including a mixing length of $\alpha_{\rm mlt}=1.8$, core overshooting of 0.2 pressure scale heights during H/He burning phases (Stancliffe et al., 2015), and semi-convection efficiency $\alpha_{sc}=0.1$ (Choi et al., 2016). Stellar wind mass loss follows the prescriptions of Marchant et al. (2016). These chosen input parameters are closely similar to those employed by Schneider et al. (2021), ensuring consistency when applying their results to determine BH progenitor masses.
@@ III.2 Effect of kick
In our default model, we assume the BH forms with zero natal kick velocity. Under this assumption, the sudden mass loss during the SN explosion invariably widens the orbital separation (from Stage 5 to Stage 6 in Figure 1), allowing the post-CE binary separations to be directly derived from the parameter spaces shown in Figures 4-6. However, the scenario where BHs receive substantial natal kicks cannot be entirely ruled out, as suggested by observational evidence from some BH binaries (e.g., Tauris et al. 1999; Repetto and Nelemans 2015; Mirabel 2017; Burrows et al. 2025; Nagarajan and El-Badry 2025; Abdulghani et al. 2025). If a BH forms with a significant kick, the random orientation of the kick can either increase or decrease the orbital separation, thereby modifying the allowable parameter space to some degree, and consequently affecting the inferred CE efficiencies. In this section, we examine how BH natal kicks influence our previous results.
@@ III.2.2 Influence of kick on BH born mass
The influence of BH natal kicks extends beyond 4U 1543 to affect the parameter spaces of SAX J1819 and GRO J1655 as well. Figure 8 displays the BH mass and companion star mass distributions corresponding to the grid points in Figures 4-6. The upper panel distinguishes between forbidden grids (gray open circles with crosses) and viable systems when BH kicks are considered (cyan circles). The middle and lower panels compare systems without BH kicks (red open circles, representing allowable grids from Figures 4 and 5) to those with kicks. Notably, BH natal kicks significantly expand the allowable parameter space. Most remarkably, for both SAX J1819 and GRO J1655, the inclusion of natal kicks allows for BH formation in the so-called mass-gap region $\sim 3-5\,M_{\odot}$ (Shao, 2022), a scenario prohibited in the zero-kick case.
@@ III.3 Progenitor Mass Dependence of Minimum $\alpha_{\rm CE}$
Figure 10: The minimum values of CE efficiency for GRO J1655 vary with the initial primary mass and the stellar radius. The left, middle and right panels correspond to the $\alpha_{\rm 0.5U}$, $\alpha_{\rm U}$ and $\alpha_{\rm H}$ cases respectively. Dashed lines indicate constant minimum CE efficiency values. Notably, when enthalpy is included (right panel), the minimum CE efficiencies show significantly less variation, typically confined to a narrow range of 1.531.5-3.
@@
We see that the minimum CE efficiency does not always increase monotonically with the mass of the BH progenitor when other parameters are held fixed. This behavior can be understood as follows. For a BH progenitor filling its Roche lobe at a given evolutionary stage, a more massive progenitor typically produces a more massive He star after CE ejection. According to Equation (5-7), a more massive He star tends to result in a wider post-BH-formation orbit. Inversely, for a specific grid point in the parameter space, a more massive progenitor would lead to a shorter post-CE orbital separation. Consequently, although a more massive BH progenitor generally has a larger envelope binding energy, the shorter post-CE separation can yield a lower estimated CE efficiency (according to Equation 1). Furthermore, the allowable parameter space varies with the initial black hole progenitor mass, which affects the derived minimum CE efficiency values. This dependence explains why the 33 $M_{\odot}$ progenitor exhibits relatively lower minimum CE efficiency compared to the 32 $M_{\odot}$ case when both systems are undergoing common-envelope evolution at similar evolutionary stages.
@@ IV Discussion
  • Extra energy sources - Nuclear energy: Podsiadlowski et al. (2010) suggested that nuclear energy could contribute to envelope ejection. In this scenario, hydrogen-rich material from the low-mass companion is injected into the helium-burning shell of the massive donor during the late CE phase, triggering a nuclear runaway that explosively ejects both the hydrogen and helium layers. The result is a close binary containing a compact object and a low-mass companion. This channel is, however, very rare and requires that the low-mass companion overflows its Roche lobe during the self-regulated spiral-in phase to transfer hydrogen-rich material deep into the donor’s helium shell (see also Ivanova, 2002; Ivanova et al., 2020). This channel involves complex mass transfer processes within the common envelope that are difficult to model with current simulations, and therefore remains a topic for future work.
@@
  • Extra energy sources - Jets: The possibility that jets launched from the accreting companion serve as an additional energy source during common envelope evolution has gained substantial support from recent theoretical, numerical, and observational studies. Numerical simulations have demonstrated that jets can remove mass before and during the CE phase, reduce drag forces, and facilitate envelope unbinding (e.g., Soker 2004, 2015; Shiber and Iaconi 2024; Weiner and Soker 2025). Furthermore, the newly proposed jetted mass-removal accretion mechanism indicates that MS stars can accrete at high rates, provided that jets remove the outer, high-entropy layers of the accreted envelope Bear and Soker 2025; Scolnic et al. 2025; Cohen et al. 2026; but see Zou et al. 2022). Observationally, jets appear to be the most robust observable ingredient of CE evolution, found in about $40\%$ more post-CE planetary nebulae than dense equatorial outflows Soker 2025. Collectively, these developments suggest that jets may be regarded as a primary candidate for the missing energy source in CE evolution. However, quantitative simulations are still needed to accurately determine the contribution of jets to envelope unbinding during the CE phase.
@@
  • Tidal effects: Tidal interactions could indirectly influence the estimated CE efficiency. If sufficiently strong, tides can significantly tighten the orbital separation as the massive star evolves into a RSG. In this scenario, the donor star may not initially fill its Roche lobe at its maximum radial extent, enabling continuous mass loss during the RSG phase as tides gradually shrink the orbit. By the onset of the CE phase (when the star finally fills its Roche lobe), a substantial portion of the envelope may have already been lost. This process is quite similar to the grazing envelope evolution (GEE) proposed by Soker (2015), in which jets remove some envelope mass before the CE phase (see also Shiber and Iaconi 2024), thereby substantially reducing the binding energy (Kotko et al., 2024). However, it is important to emphasize that current tidal theory still contains significant uncertainties, particularly regarding the efficiency of orbital decay (e.g., Meibom and Mathieu, 2005; Ogilvie, 2014; Nie et al., 2017).
@@
  • Alternative CE description: Recent work by Hirai and Mandel (2022) has proposed a novel two-stage CE ejection mechanism for massive binaries, motivated by the stellar structure of RSGs that typically exhibit a substantial radiative layer between their dense He cores and outer convective envelopes. In this model, CE evolution proceeds through distinct phases, i.e., rapid inspiral through the convective envelope, followed by stable mass transfer from the radiative intershell region (see also Marchant et al. 2021; Gallegos-Garcia et al. 2023; Nie et al. 2025). Given that the later phase involves stable mass transfer between the core (radiative intershell region) and the companion, the final post-CE separation is determined by the angular momentum loss mechanism and the specific binary parameters prevailing during the mass transfer epoch. However, recent simulations by Cohen and Soker (2023) suggested that the inner boundary of the envelope convective zone moves deep into the initial radiative zone, highlighting the uncertainties inherent in this formalism.

### 26 Apr 2026 v2

@@ Abstract
The massive binary common envelope (CE) phase plays a pivotal role in the formation of close black hole/neutron star (BH/NS) binaries, yet significant uncertainties remain in our understanding of this process. In this study, we aim to constrain the massive binary CE phase by systematically reconstructing three observed BH X-ray binaries (BHXBs): GRO J1655-40, SAX J1819.3-2525, and 4U 1543-47. Through comprehensive binary evolution simulations and parametric supernova modeling, we establish lower limits for the CE efficiency parameters under different energy considerations within the standard energy formalism. Specifically, we derive minimum values for three cases: $\alpha_{\rm 0.5U}$ and $\alpha_{\rm U}$ representing CE efficiencies with half and all of the internal energy contributing to the envelope ejection, respectively, and $\alpha_{\rm H}$ accounting for the envelope’s enthalpy. Our analysis reveals that the self-consistent formation of these three BHXBs requires CE efficiency parameters satisfying: $\alpha_{\rm 0.5U}\gtrsim 6.7$, $\alpha_{\rm U}\gtrsim 4.2$ and $\alpha_{\rm H}\gtrsim 1.7$. Notably, we find no viable solutions with CE efficiency values below unity, even when considering the most extreme scenarios in which the envelope binding energy is significantly reduced through enthalpy inclusion. Our results strongly imply that either additional energy sources are required, or the formalism itself must be revised. Furthermore, we quantitatively assess the impact of BH natal kicks on our results. A key finding is that 4U 1543-47’s formation requires substantial natal kicks ($\gtrsim 50\;\rm km/s$), as lower kick velocities are incompatible with isolated binary evolution.
@@
software: MESA (v12115; Paxton et al. 2011, 2013, 2015, 2018, 2019; Jermyn et al. 2023)
@@ I Introduction
Although significant effort has been devoted to modeling CE evolution through numerical simulations spanning 1D to 3D approaches (e.g., Soker, 2016; Ohlmann et al., 2016; Ivanova and Nandez, 2016; Kruckow et al., 2016; Iaconi et al., 2017; Shiber and Soker, 2018; Chamandy et al., 2018; Fragos et al., 2019; Shiber et al., 2019; Law-Smith et al., 2020; Grichener et al., 2021; Moreno et al., 2022; Lau et al., 2022; Bronner et al., 2024; Gagnier and Pejcha, 2025; Vetter et al., 2025), the astrophysical community has largely adopted a phenomenological description based on energy conservation as a practical framework for modeling the CE ejection (Paczynski, 1976; Webbink, 1984; Livio and Soker, 1988). In this energy formalism [^1] In the remainder of this paper, the CE energy formalism refers to the standard version used in binary population synthesis studies. It therefore does not consider additional processes or energy sources that may operate during the CE phase, including but not limited to jets (e.g., Soker 2004, 2015, 2016; Shiber et al. 2019), magnetic fields (e.g., Ohlmann et al. 2016; García-Segura et al. 2020) and nuclear energy (Podsiadlowski et al., 2010)., the system’s orbital energy release is assumed to overcome the envelope’s binding energy. This approach offers significant advantages: its simplicity and computational efficiency make it particularly suitable for population synthesis studies (see recent reviews by Han et al. 2020; Chen et al. 2024).
@@ II.2 Formation scenario of BHXBs
Figure 1: The typical evolutionary scenario for BH I/LMXBs. Abbreviations are as follows: ZAMS–zero-age main sequence, MT–mass transfer, CE–common envelope, He–helium, MS–main sequence, SN–supernova, BH–black hole, I/LMXB–intermediate-/low-mass X-ray binary. At stage 4, the dashed contour denotes the Roche lobe of the secondary star. Note that the successful ejection of the CE requires the secondary’s radius to be within its Roche lobe.
@@ II.3 The technical route
where $P$ and $\rho$ represent the pressure and density of each mass shell. The quantity $U+P/\rho$ is generally known as enthalpyenthalpy $(H)$. It should be noted that enthalpy consideration does not introduce new energy sources, but rather redistributes the deposited heat during the spiral-in phase (Ivanova and Chaichenets, 2011; Ivanova et al., 2013). Following Ivanova (2011), we define the core-envelope boundary as the point with the maximum compression $P/\rho$ inside the hydrogen burning shell (characterized by the peak values of the local sonic velocity; see also Marchant et al. 2021; Vigna-Gómez et al. 2022), a choice motivated by the need to prevent the donor remnant from expanding immediately after CE ejection.
Figure 2: Binding energy calculation example for a $34\,M_{\odot}$ primary star. Panel (a): The evolutionary track in the HR diagram with the red cross indicating the specific evolutionary stage analyzed. The thick green line indicates that the star contains a He core. Panel (b): Radial profiles of hydrogen mass fraction ($X_{H}$) and the sonic velocity $C_{\rm s}\equiv\sqrt{P/\rho}$ at this stage, where open circles mark the core-envelope boundary. Panel (c): The evolution of the primary envelope energy components (gravitational energy $E_{\rm gr}$, internal energy $E_{\rm U}$, and enthalpy term $E_{\rm H}$), as functions of stellar radius at each specific evolutionary stage (covering the post-MS stages). Panel (d): Similar to panel (c), but for the integrated envelope binding energy vs. radius for the three cases.
@@
From Stage 5 to Stage 4, the orbit evolution is governed by mass loss through the stripped He star’s stellar wind, which we consider as a fast (several $1000\;\rm km/s$; Vink 2022) and isotropic spherically-symmetric wind that does not interact with the companion star. The assumptions are well-justified for Wolf-Rayet stars where the wind velocity significantly exceeds the typical orbital velocity ($\sim 100\;\rm km/s$), making wind accretion negligible (El Mellah et al., 2019; Li et al., 2026). The change in binary separation due to the wind mass loss then is given by (Postnov and Yungelson, 2014)
@@
For each set of binary parameters at Stage 6 that successfully reproduces the observed BHXB properties, we can systematically reconstruct the corresponding post-CE binary configuration at Stage 4. Using Equation (1) together with the binding energy $E_{\rm bind}$ calculated for a given BH progenitor across different evolutionary stages, we determine the required CE efficiency parameter $\alpha_{\rm CE}$ for each viable solution.
@@ II.5 The input parameters
We perform all stellar evolution calculations using the state-of-the-art stellar evolutionary code Modules for Experiments in Stellar Astrophysics (MESA; version 12115; Paxton et al. 2011, 2013, 2015, 2018, 2019; Jermyn et al. 2023), adopting solar metallicity (Z=0.0142) with initial He fraction Y=0.2703 (Asplund et al., 2009) for non-rotating massive star models (as the strong stellar winds in massive stars cause a marked spin-down; e.g., Nathaniel et al. 2025). Our simulations incorporate Type II opacities (Iglesias and Rogers, 1996) and the approx21_\_cr60_\_plus_\_co56.net nuclear reaction network (Ferguson et al., 2005), with convection parameters including a mixing length of $\alpha_{\rm mlt}=1.8$, core overshooting of 0.2 pressure scale heights during H/He burning phases (Stancliffe et al., 2015), and semi-convection efficiency $\alpha_{\rm sc}=0.1$ (Choi et al., 2016). Stellar wind mass loss follows the prescriptions of Marchant et al. (2016). These chosen input parameters are closely similar to those employed by Schneider et al. (2021), ensuring consistency when applying their results to determine BH progenitor masses.
@@ III.2 Effect of kick
In our default model, we assume the BH forms with zero natal kick velocity. Under this assumption, the sudden mass loss during the SN explosion invariably widens the orbital separation (from Stage 5 to Stage 6 in Figure 1), allowing the post-CE binary separations to be directly derived from the parameter spaces shown in Figures 4-6. However, the scenario where BHs receive substantial natal kicks cannot be entirely ruled out, as suggested by observational evidence from some BH binaries (e.g., Tauris et al. 1999; Repetto and Nelemans 2015; Mirabel 2017; Zhao et al. 2023; Dashwood Brown et al. 2024; Burrows et al. 2025; Nagarajan and El-Badry 2025; Abdulghani et al. 2025). If a BH forms with a significant kick, the random orientation of the kick can either increase or decrease the orbital separation, thereby modifying the allowable parameter space to some degree, and consequently affecting the inferred CE efficiencies. In this section, we examine how BH natal kicks influence our previous results.
@@ III.2.2 Influence of kick on BH born mass
The influence of BH natal kicks extends beyond 4U 1543 to affect the parameter spaces of SAX J1819 and GRO J1655 as well. Figure 8 displays the BH mass and companion star mass distributions corresponding to the grid points in Figures 4-6. The upper panel distinguishes between forbidden grids (gray open circles with crosses) and viable systems when BH kicks are considered (cyan circles). The middle and lower panels compare systems without BH kicks (red open circles, representing allowable grids from Figures 4 and 5) to those with kicks. Notably, BH natal kicks significantly expand the allowable parameter space. Most remarkably, for both SAX J1819 and GRO J1655, the inclusion of natal kicks allows for BH formation in the so-called mass-gap region $\sim 3-5\,M_{\odot}$ (Shao, 2022; Wang et al., 2024), a scenario prohibited in the zero-kick case.
@@ III.3 Progenitor Mass Dependence of Minimum $\alpha_{\rm CE}$
Figure 10: The minimum values of CE efficiency for GRO J1655 vary with the initial primary mass and the stellar radius. The left, middle and right panels correspond to the $\alpha_{\rm 0.5U}$, $\alpha_{\rm U}$ and $\alpha_{\rm H}$ cases respectively. The shaded gray region indicates CE efficiencies above the upper limit of the colorbar. Dashed lines indicate constant minimum CE efficiency values. Notably, when enthalpy is included (right panel), the minimum CE efficiencies show significantly less variation, typically confined to a narrow range of 1.531.5-3.
@@
We see that the minimum CE efficiency does not always increase monotonically with the mass of the BH progenitor when other parameters are held fixed. This behavior can be understood as follows. For a BH progenitor filling its Roche lobe at a given evolutionary stage, a more massive progenitor typically produces a more massive He star after CE ejection. According to Equation (5-7), a more massive He star tends to result in a wider post-BH-formation orbit. Inversely, for a specific grid point in the parameter space, a more massive progenitor would lead to a shorter post-CE orbital separation. Consequently, although a more massive BH progenitor generally has a larger envelope binding energy, the shorter post-CE separation can yield a lower estimated CE efficiency (according to Equation 1). Furthermore, the allowable parameter space varies with the initial black hole progenitor mass, which affects the derived minimum CE efficiency values. This dependence explains why the 33 $M_{\odot}$ progenitor exhibits relatively lower minimum CE efficiency compared to the 32 $M_{\odot}$ case when both systems are undergoing CE evolution at similar evolutionary stages.
@@ IV Discussion
  • Core-envelope boundary: One of the largest uncertainties in calculating the envelope binding energy lies in the definition of the core-envelope boundary. In this work, we adopt a physically motivated prescription, where the boundary is determined by the maximum compression point (Ivanova, 2011). Under this assumption, the core-envelope boundary typically corresponds to a hydrogen mass fraction of $X_{\rm H}\approx 0.1$$0.3$. This adopted boundary is located farther out (i.e., closer to the stellar surface) compared to the commonly used criterion in binary population synthesis studies, which often take $X_{\rm H}=0.1$ as the core-envelope boundary (e.g., Kruckow et al. 2016, 2018). Moving the boundary outward would undoubtedly reduce the envelope binding energy and thus lower the inferred $\alpha_{\rm CE}$ values. However, a more outward boundary also implies that more hydrogen remains in the envelope, which may lead to significant re-expansion of the stripped core after the CE phase (e.g., De Marco et al. 2011; Marchant et al. 2021; Vigna-Gómez et al. 2022). Recent simulations by Vigna-Gómez et al. (2022) examined the stellar response following the removal of the hydrogen-rich envelope. They found that the boundary allowing the star to avoid rapid re-expansion lies above the maximum compression point. Nevertheless, their simulations focused on stars with masses below $25\,M_{\odot}$, while our work primarily targets more massive stars ($\geq 32\,M_{\odot}$). This mass difference may result in different post-stripping behaviour. A key distinction is that for stars below $25\,M_{\odot}$, the maximum compression point is generally located at $X_{\rm H}<0.1$ (Vigna-Gómez et al., 2022). In contrast, for stars with masses $\gtrsim 32\,M_{\odot}$, we find that the maximum compression point falls in the range $0.1<X_{\rm H}<0.3$. Therefore, both the appropriate definition of the core-envelope boundary and the post-CE evolution of the stripped core remain open questions that warrant further investigation.
@@
  • Extra energy sources - Nuclear energy: Podsiadlowski et al. (2010) suggested that nuclear energy could contribute to envelope ejection. In this scenario, hydrogen-rich material from the low-mass companion is injected into the helium-burning shell of the massive donor during the late CE phase, triggering a nuclear runaway that explosively ejects both the hydrogen and helium layers. The result is a close binary containing a compact object and a low-mass companion. This channel is, however, very rare and requires that the low-mass companion overflows its Roche lobe during the self-regulated spiral-in phase to transfer hydrogen-rich material deep into the donor’s helium shell (see also Ivanova, 2002; Ivanova et al., 2020). This channel involves complex mass transfer processes within the CE that are difficult to model with current simulations, and therefore remains a topic for future work.
@@
  • Extra energy sources - Jets: The possibility that jets launched from the accreting companion serve as an additional energy source during CE evolution has gained substantial support from recent theoretical, numerical, and observational studies. Numerical simulations have demonstrated that jets can remove mass before and during the CE phase, reduce drag forces, and facilitate envelope unbinding (e.g., Soker 2004, 2015; Shiber and Iaconi 2024; Weiner and Soker 2025). Furthermore, the newly proposed jetted mass-removal accretion mechanism indicates that MS stars can accrete at high rates, provided that jets remove the outer, high-entropy layers of the accreted envelope (Bear and Soker 2025; Scolnic et al. 2025; Cohen et al. 2026; but see Zou et al. 2022). Observationally, jets appear to be the most robust observable ingredient of CE evolution, found in about $40\%$ more post-CE planetary nebulae than dense equatorial outflows (Soker, 2025). Collectively, these developments suggest that jets may be regarded as a primary candidate for the missing energy source in CE evolution. However, quantitative simulations are still needed to accurately determine the contribution of jets to envelope unbinding during the CE phase.
@@
  • Tidal effects: Tidal interactions could indirectly influence the estimated CE efficiency. If sufficiently strong, tides can significantly tighten the orbital separation as the massive star evolves into a RSG. In this scenario, the donor star may not initially fill its Roche lobe at its maximum radial extent, enabling continuous mass loss during the RSG phase as tides gradually shrink the orbit. By the onset of the CE phase (when the star finally fills its Roche lobe), a substantial portion of the envelope may have already been lost. This process is quite similar to the grazing envelope evolution proposed by Soker (2015), in which jets remove some envelope mass before the CE phase (see also Shiber and Iaconi 2024), thereby substantially reducing the binding energy (Kotko et al., 2024). However, it is important to emphasize that current tidal theory still contains significant uncertainties, particularly regarding the efficiency of orbital decay (e.g., Meibom and Mathieu, 2005; Ogilvie, 2014; Nie et al., 2017).
@@
  • Alternative CE description: Recent work by Hirai and Mandel (2022) has proposed a novel two-stage CE ejection mechanism for massive binaries, motivated by the stellar structure of RSGs that typically exhibit a substantial radiative layer between their dense He cores and outer convective envelopes. In this model, CE evolution proceeds through distinct phases, i.e., rapid inspiral through the convective envelope, followed by stable mass transfer from the radiative intershell region (see also Marchant et al. 2021; Gallegos-Garcia et al. 2023; Nie et al. 2025). Given that the later phase involves stable mass transfer between the core (radiative intershell region) and the companion, the final post-CE separation is determined by the angular momentum loss mechanism and the specific binary parameters prevailing during the mass transfer epoch. However, recent simulations by Cohen and Soker (2023) suggested that the inner boundary of the envelope convective zone moves deep into the initial radiative zone, highlighting the uncertainties inherent in this formalism. Apart from the energy-based CE prescription, the $\gamma$-mechanism (e.g., Nelemans et al. 2000; Nelemans and Tout 2005; Di Stefano et al. 2023) treats the CE phase via angular momentum conservation. Although the $\gamma$-mechanism appears to reproduce some observed binary systems better than the energy formalism (e.g., Nelemans et al., 2025), the underlying physical picture remains unclear.